# RiemannFormer: A Framework for Attention in Curved Spaces

### v1

@@ 2.1 Locality Focusing
Unlike CNNs with locality and translational equivariance, vanilla transformers are data-hungry due to their minimal inductive biases. This flexibility makes them powerful on large, diverse datasets but fragile on limited data, where domain-specific priors are crucial for generalization. To enhance the local inductive bias of transformers, we introduce an explicit mechanism to highlight a neighborhood by attenuating the remote values.
Assume that the tokens are placed in a flat space and the explicit positional embedding are also absent, is it possible to introduce an additional mechanism to address the lack of awareness of sequence order in self-attention module?
To clarify, rather than considering the values $\mathbf{V}$ in a curved space, we aggregate them in the parameter space $\mathbb{R}^{D}$. Given a position $m$, we introduce a decay factor for the values $\mathbf{V}$ to attenuate the influence of remote positions. Specifically, we introduce factors $\lambda_{i}$ to scale the values $\mathbf{v}_{i}$ and update the attention function $\sum_{i}s_{i}(\lambda_{i}\mathbf{v}_{i})$. We compute the output vector as:
@@
$\mathbf{v}_{m}^{\prime}(\mathbf{q}_{m})=\text{Attention}(\mathbf{q}_{m},\mathbf{K},\mathbf{\Lambda},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{q}_{m}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\right)(\mathbf{\Lambda}\mathbf{V})=\mathbf{s}(\mathbf{\Lambda}\mathbf{V}),$ (19)
@@
where $\mathbf{q}_{m},\;\mathbf{s}\in\mathbb{R}^{1\times D}$, $\mathbf{K},\;\mathbf{V}\in\mathbb{R}^{L\times D}$, and a diagonal matrix $\mathbf{\Lambda}\in\mathbb{R}^{L\times L}$. The formula can be expanded as follows,
$\mathbf{v}_{m}^{\prime}(\mathbf{q}_{m})$ $=\begin{bmatrix}s_{m1}&s_{m2}&\cdots&s_{mL}\end{bmatrix}\begin{bmatrix}\lambda_{m1}&&&\\ &\lambda_{m2}&&\\ &&\ddots&\\ &&&\lambda_{mL}\end{bmatrix}\begin{bmatrix}\mathbf{v}_{1}\\[3.6pt] \mathbf{v}_{2}\\ \vdots\\ \mathbf{v}_{L}\end{bmatrix}$ (20)
$=\begin{bmatrix}s_{m1}&s_{m2}&\cdots&s_{mL}\end{bmatrix}\begin{bmatrix}\lambda_{m1}\mathbf{v}_{1}\\[3.6pt] \lambda_{m2}\mathbf{v}_{2}\\ \vdots\\ \lambda_{mL}\mathbf{v}_{L}\end{bmatrix}.$
Here, we discuss $2D$ image data, although this mechanism is equally applicable to $nD$ sequence data. The output value $\mathbf{v}^{\prime}_{m}$ can also be interpreted from the perspective of a bilateral filter [16] in the context of image processing,
$\mathbf{v}^{\prime}_{m}=\sum_{n=1}^{L}{G}_{s}(\mathbf{q}_{m},\mathbf{k}_{n})G_{\sigma}(\mathbf{p}_{m},\mathbf{p}_{n})\mathbf{v}_{n},$ (21)
where $G_{s}(\cdot,\cdot)$ indicates the scaled dot-product of the query and key vectors followed a softmax function, $\mathbf{p}_{m}$ and $\mathbf{p}_{n}$ are the $2D$ positions of image patches on the image grid, $G_{\sigma}(\cdot,\cdot)$ is an attenuation factor calculated based on the norm of the difference between $\mathbf{x}$ and $\mathbf{y}$, such as $G_{\sigma}(\mathbf{x},\mathbf{y})=e^{-\frac{\|\mathbf{x}-\mathbf{y}\|^{2}_{A}}{2\sigma^{2}}}$. $\mathbf{A}$ is a symmetric positive-definite matrix such that $\|\mathbf{v}\|_{\bf A}^{2}=\mathbf{v}^{\top}\mathbf{A}\mathbf{v}$, it is learnable and is set to the identity matrix by default. $\sigma$ is a learnable parameter that can be assigned a unique value for each position or shared as one parameter for all positions in an attention head. From the perspective of the bilateral filter, $G_{s}(\cdot,\cdot)$ controls intensity similarity and $G_{\sigma}(\cdot,\cdot)$ controls spatial attenuating.
In the implementation, the attenuation factors corresponding to one position are reshaped into a row, thus the ones of all positions are packed together into a $L\times L$ attenuation matrix. For all tokens, we can first perform element-level multiplication on the score matrix and the attenuation matrix $\mathbf{\Omega}$, and then perform matrix multiplication on the result matrix and $\mathbf{V}$. We have
@@
$(\mathbf{S}\odot\mathbf{\Omega})\cdot\mathbf{V}=\left(\begin{bmatrix}s_{11}&s_{12}&\cdots&s_{1L}\\ s_{21}&s_{22}&\cdots&s_{2L}\\ \vdots&\vdots&\ddots&\vdots\\ s_{L1}&s_{L2}&\cdots&s_{LL}\end{bmatrix}\odot\begin{bmatrix}\lambda_{11}&\lambda_{12}&\cdots&\lambda_{1L}\\ \lambda_{21}&\lambda_{22}&\cdots&\lambda_{2L}\\ \vdots&\vdots&\ddots&\vdots\\ \lambda_{L1}&\lambda_{L2}&\cdots&\lambda_{LL}\end{bmatrix}\right)\cdot\mathbf{V}.$ (22)
@@
Such a simple mechanism is introduced to improve the capacity of transformers to capture and learn the inductive bias of locality especially on small datasets. Our experimental results reveal that incorporating such an explicit local focusing mechanism substantially improves the performance of transformers in visual tasks.
@@ 3 Experiments
In the context of image data, there are two axes to consider. Although these axes can be addressed jointly [14] or the image patches are directly flatten into a sequence, we are interested in exploring the effects of processing them separately along the $x$ and $y$ axes. This strategy also facilitates a comparative studies with previous work.
@@
In predefined transformations, certain operations (such as rotation $\breve{\bf R}(\theta)$ and reflection $\hat{\bf R}(\theta)$) involve the configuration of angles. An angle schedule is a function that sends an integer to an angle involved in the linear transformation. The choice of the angle schedule can also impact the performance of the tangent mapping. In our experiments, we observed that the following straightforward schedule yields favorable results: $f(t):t\mapsto t\theta,\;t\in\{1,2,\dots\}$, where $\theta$ is a learnable parameter or a predefined angle. Exploring more effective angle schedules is also a direction worth investigating. In the current experiment of this work, we use the predefined configuration in Eq. (10). We set the hyperparameter $\alpha=0.1$ in $s=\frac{\exp(w)}{\exp(w)+\alpha}$ in these experiments.
@@ 4 Conclusion
$\mathbf{S}=\left[\begin{matrix}s_{1}&0\\[8.09998pt] 0&s_{2}\end{matrix}\right]$
@@
$=\mathbf{q}_{m}^{\top}\left[\breve{\bf R}(-m\theta)\cdot\mathbf{S}^{\frac{m+n}{2}}\cdot\breve{\bf R}(n\theta)\right]\mathbf{k}_{n}$
@@
$\langle s^{\frac{m}{2}}\breve{\bf R}(m\theta)\mathbf{q}_{m},\;s^{\frac{n}{2}}\breve{\bf R}(n\theta)\mathbf{k}_{n}\rangle_{\mathbf{I}}=\langle\mathbf{T}_{m}^{-1}\mathbf{q}_{m},\mathbf{T}_{n}^{-1}\mathbf{k}_{n}\rangle_{\mathbf{I}}.$

### v2

@@ A Concrete Realization: The Locality Focusing Mechanism
Unlike CNNs, which are endowed with locality and translational equivariance, vanilla transformers are data-hungry due to their minimal inductive biases. This flexibility makes them powerful on large, diverse datasets but can be a limitation on limited data, where domain-specific priors are crucial for generalization. Our Locality Focusing mechanism directly addresses this by introducing an explicit neighborhood bias, attenuating the influence of remote values during aggregation.
Consistent with the trivial bundle structure that places value vectors in a global Euclidean space $\mathbb{R}^{D}$, this modulation is performed directly on the values prior to their final weighting. While attention scores implicitly reflect semantic relevance, Locality Focusing enforces a strong spatial bias. Specifically, for a given query position $m$, we modulate each value vector $\mathbf{v}_{i}$ by a position-relative decay factor, $\lambda_{m,i}$. This factor is designed to be a monotonic function of the distance between positions $m$ and $i$. The final output vector for position $m$ is then computed as the weighted sum of these modulated values, where $s_{m,i}$ are the softmax-normalized attention scores:
$\mathbf{o}_{m}=\sum_{i}s_{m,i}\left(\lambda_{m,i}\mathbf{v}_{i}\right).$
This approach effectively disentangles the "what" (semantic relevance from attention) from the "where" (spatial locality from the $\lambda$ factors), providing a more robust and controllable inductive bias. In matrix form, the output for a single query $\mathbf{q}_{m}$ is:
@@
$\mathbf{o}_{m}=\text{Attention}(\mathbf{q}_{m},\mathbf{K},\mathbf{\Lambda}_{m},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{q}_{m}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\right)(\mathbf{\Lambda}_{m}\mathbf{V})=\mathbf{s}_{m}(\mathbf{\Lambda}_{m}\mathbf{V}),$ (12)
@@
where $\mathbf{s}_{m}\in\mathbb{R}^{1\times L}$, $\mathbf{K},\mathbf{V}\in\mathbb{R}^{L\times D}$, and $\mathbf{\Lambda}_{m}\in\mathbb{R}^{L\times L}$ is a diagonal matrix containing the factors $\{\lambda_{m,i}\}_{i=1}^{L}$. This expands to:
$\mathbf{o}_{m}$ $=\begin{bmatrix}s_{m,1}&s_{m,2}&\cdots&s_{m,L}\end{bmatrix}\begin{bmatrix}\lambda_{m,1}&&&\\ &\lambda_{m,2}&&\\ &&\ddots&\\ &&&\lambda_{m,L}\end{bmatrix}\begin{bmatrix}\mathbf{v}_{1}\\[3.6pt] \mathbf{v}_{2}\\ \vdots\\ \mathbf{v}_{L}\end{bmatrix}$ (13)
$=\begin{bmatrix}s_{m,1}&s_{m,2}&\cdots&s_{m,L}\end{bmatrix}\begin{bmatrix}\lambda_{m,1}\mathbf{v}_{1}\\[3.6pt] \lambda_{m,2}\mathbf{v}_{2}\\ \vdots\\ \lambda_{m,L}\mathbf{v}_{L}\end{bmatrix}$.
This mechanism, while presented here for 2D image data, is equally applicable to n-dimensional sequences. Its formulation is strongly reminiscent of the bilateral filter from image processing, which smooths images while preserving edges:
$\mathbf{o}_{m}=\sum_{n=1}^{L}\underbrace{G_{s}(\mathbf{q}_{m},\mathbf{k}_{n})}_{\text{Intensity Similarity}}\underbrace{G_{\sigma}(\mathbf{d}_{m},\mathbf{d}_{n})}_{\text{Spatial Proximity}}\mathbf{v}_{n},$ (14)
Here, the attention term $G_{s}(\cdot,\cdot)$ plays the role of the intensity/range kernel, measuring similarity in feature space, while our attenuation factor $G_{\sigma}(\cdot,\cdot)$ acts as the spatial/domain kernel. A common choice for $G_{\sigma}$ is a Gaussian function of the distance between positions $\mathbf{d}_{m}$ and $\mathbf{d}_{n}$, $G_{\sigma}(\mathbf{x},\mathbf{y})=e^{-\frac{\|\mathbf{x}-\mathbf{y}\|^{2}_{A}}{2\sigma^{2}}}$, where $\mathbf{A}$ and $\sigma$ can be learnable.
For efficient implementation across all tokens, the attenuation factors for all query positions can be collected into a single $L\times L$ matrix $\mathbf{\Omega}$, where $\mathbf{\Omega}_{m,i}=\lambda_{m,i}$. The full attention output is then computed as:
@@
$\mathbf{O}=(\mathbf{S}\odot\mathbf{\Omega})\mathbf{V}=\left(\begin{bmatrix}s_{1,1}&s_{1,2}&\cdots&s_{1,L}\\ s_{2,1}&s_{2,2}&\cdots&s_{2,L}\\ \vdots&\vdots&\ddots&\vdots\\ s_{L,1}&s_{L,2}&\cdots&s_{L,L}\end{bmatrix}\odot\begin{bmatrix}\lambda_{1,1}&\lambda_{1,2}&\cdots&\lambda_{1,L}\\ \lambda_{2,1}&\lambda_{2,2}&\cdots&\lambda_{2,L}\\ \vdots&\vdots&\ddots&\vdots\\ \lambda_{L,1}&\lambda_{L,2}&\cdots&\lambda_{L,L}\end{bmatrix}\right)\cdot\mathbf{V}.$ (15)
@@
This simple mechanism significantly improves the capacity of transformers to learn the inductive bias of locality, especially on smaller datasets. As our experiments reveal, incorporating such an explicit local focusing mechanism substantially improves the performance of transformers in visual tasks.
@@ 3 Experiments
In the context of image data, there are two axes to consider. Although these axes can be addressed jointly [heo2024ropevit] or the image patches are directly flatten into a sequence, we are interested in exploring the effects of processing them separately along the $x$ and $y$ axes. This strategy also facilitates a comparative studies with previous work.
@@
In predefined transformations, rotation operations in subspaces involve the configuration of angles. An angle schedule is a function that sends an integer to an angle involved in the linear transformation. The choice of the angle schedule can also impact the performance of the tangent mapping. In our experiments, we observed that the following straightforward schedule yields favorable results: $f(t):t\mapsto t\theta,\;t\in\{1,2,\dots\}$, where $\theta$ is a learnable parameter or a predefined angle. Exploring more effective angle schedules is also a direction worth investigating. In the current experiment of this work, we use 1D PaTE-SO(2) with the Conformal Exponential Scaling scheme ($\mathbf{M}_{i}=s^{i}\mathbf{I}$), and we set the hyperparameter $\alpha=0.1$ in $s=\frac{\exp(w)}{\exp(w)+\alpha}$ in these experiments.
@@ 4 Conclusion
$\mathbf{S}=\left[\begin{matrix}s_{1}&0\\[8.09998pt] 0&s_{2}\end{matrix}\right],$
@@
$=\mathbf{q}_{m}^{\top}\left[\breve{\bf R}(-m\theta)\cdot\mathbf{S}^{\frac{m+n}{2}}\cdot\breve{\bf R}(n\theta)\right]\mathbf{k}_{n}.$
@@
$\langle\mathbf{S}^{\frac{m}{2}}\breve{\bf R}(m\theta)\mathbf{q}_{m},\;\mathbf{S}^{\frac{n}{2}}\breve{\bf R}(n\theta)\mathbf{k}_{n}\rangle_{\mathbf{I}}=\langle\mathbf{T}_{m}^{-1}\mathbf{q}_{m},\mathbf{T}_{n}^{-1}\mathbf{k}_{n}\rangle_{\mathbf{I}}.$